40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 17 20:22:07 2012, Jamie, Update, General, Vent plan: WE VENT TOMORROW 
    Reply  Mon Dec 17 21:33:45 2012, Manasa, Update, General, Vent plan: WE VENT TOMORROW 
    Reply  Tue Dec 18 07:03:40 2012, Manasa, Update, General, Vent plan: WE VENT TOMORROW 
    Reply  Tue Dec 18 10:43:20 2012, Steve, Update, VAC, vent has completed vent74.pngbvent.pngbefVent.png
       Reply  Wed Dec 19 11:53:23 2012, Steve, Update, VAC, AC removed workingatatm.png
       Reply  Thu Jan 24 14:40:09 2013, Steve, Update, VAC, RGA backgroung Screenshot.png
       Reply  Mon Feb 11 16:16:27 2013, Steve, Update, VAC, 55 days at atmoshere atm55d.png
Message ID: 7843     Entry time: Mon Dec 17 21:33:45 2012     In reply to: 7842
Author: Manasa 
Type: Update 
Category: General 
Subject: Vent plan: WE VENT TOMORROW 

Quote:

Manasa is lowering the input power and preparing the mode cleaner.

I aligned the PSL angle and position QPDs...then attenuated the input laser power using HWP-PBS-HWP combo from 1.28W to 100mW following instructions from elog 6892 and elog 7299.

To enable MC locking, I replaced the 10% BS before the MC_REFL PD with Y1 mirror. I tweaked the steering mirrors at the PSL table a teeny tiny bit to enable MC locking. MC is now locked at low power with 1.0 transmission.

ELOG V3.1.3-